Additional Exercise 13.123 The rate at which crickets chirp depends on the ambient temperature, because crickets are cold bloWhat is the answer to part (b), not in scientific notation?
Answer #1

a)T = 35,number of chirps= 35-4 = 31

b)My obtained answer is quite close to your answer,it should be on the part of teacher to consider an error margin of 10% atleast as graphs could differ when made by hand as you cannot estimate the least square curve properly.

T(°C) Chirps % (K) Gc-axis) 3.418 10-3 Incherps (y-axis) 2.77 3.35x10-3 3.05 3.26 30 3. 30x 10-3 26 35 3.43 31 3.24 x 10-3 toc)T= 45 C

Number of chirps = 45-4=41